# Bias in physics peer recognition does not explain gaps in perceived recognition

### v3

@@ Abstract
Gaining recognition as a physics person from peers is an important contributor to undergraduate students’ physics identity and their success in physics courses. Previous research has separately demonstrated that women perceive less recognition from peers than men in their physics courses (perceived recognition) and that women receive fewer nominations from their peers as strong in their physics course than men (received recognition). The relationship between perceived and received peer recognition between men and women, however, is not well understood. Here we test three plausible models for this relationship. We conduct a large-scale, quantitative study of over 1,600 students enrolled in introductory physics courses at eight different institutions. We directly compare student gender, perceived recognition, and received recognition, controlling for other student demographics and course-level variability. Results show with high precision that, for students receiving the same amount of recognition, and having the same race or ethnicity, academic year, and major, women report significantly lower perceived recognition than men. These findings offer important implications for the design of effective instructional interventions.
The under-representation of women and people of color in undergraduate physics courses and, subsequently, in physics careers is well-documented [1, 2, 3, 4]. Broadening the population of physicists would give voice to new and innovative ideas that are critical to scientific advancements [5, 6, 7]. While many structural barriers to entry and retention in physics exist for women in particular, one significant challenge is that women may both perceive [8, 9, 10, 11, 12] and receive [13, 14, 15, 16, 17] less recognition from their physics peers about their abilities as physicists than men. Such recognition from peers directly correlates with students’ development of their physics identity (i.e., the extent to which they believe they are a “physics person” [18]) and their persistence in physics courses and careers [19, 9, 10, 8, 20].
Existing research in physics education, however, has only examined gender differences in students’ perceived and received peer recognition separately [13, 14, 15, 16, 17, 8, 9, 10, 11, 12]. Understanding the nuance of how students internalize the peer recognition they actually receive is crucial for identifying effective instructional interventions aimed at mitigating recognition gaps between men and women. In this study, therefore, we directly compare the extent to which students feel recognized by their peers (perceived recognition) to the number of nominations they receive from their peers for being strong in their physics course (which we assume is a proxy for received recognition), as indicated on surveys.
@@
We directly test three possible quantitative relationships between perceived and received peer recognition for men and women (Fig. 1). The first relationship proposes that both men and women internalize their peer recognition similarly. Under this relationship, men only perceive more peer recognition than women [8, 9, 10, 11, 12] because they receive more peer recognition than women (Fig. 1a), as suggested by prior work [13, 15, 16, 17]. If this relationship best describes our data and our assumption that student nominations of strong peers made on a survey are an appropriate proxy for received recognition is true, then meaningful instructional interventions should target received recognition. For example, instructors may try to refine students’ criteria through which they recognize their physics peers or diversify students’ peer networks. Quantitatively, such an intervention would aim to shift the horizontal ranges of men’s and women’s lines in Fig. 1a to be comparable.
@@
The second relationship, in contrast, is that men and women internalize their peer recognition differently. Under this relationship, men perceive more peer recognition than women [8, 9, 10, 11, 12] because they have higher perceptions of their peer recognition than women receiving the same amount of recognition (though men may also receive more peer recognition than women; Fig. 1b). If this relationship best describes our data (and our assumption that student nominations of strong peers made on a survey are an appropriate proxy for received recognition is true), instructional interventions should target perceived recognition to ensure that all students appropriately internalize recognition from their physics peers. Quantitatively, such an intervention would aim to vertically shift men’s and women’s lines in Fig. 1b to be overlapping (a supplemental intervention aimed at shifting the horizontal ranges of men’s and women’s lines in Fig. 1b may also be appropriate if there is also a gender bias in received recognition).
@@
The third relationship is that men and women internalize their peer recognition at different rates. Under this relationship, the difference between men’s and women’s perceived recognition is different at different amounts of received recognition (and men may also receive more peer recognition than women; Fig. 1c). If this relationship best describes our data (and our assumption that student nominations of strong peers made on a survey are an appropriate proxy for received recognition is true), the appropriate instructional intervention would depend on the particular relationship. An intervention should target perceived recognition (i.e., vertically shifting men’s and women’s lines in Fig. 1c), for example, if the difference between men’s and women’s perceived recognition increases for higher amounts of received recognition, as shown in Fig. 1c. On the other hand, an intervention targeting received recognition (i.e., horizontally shifting men’s and women’s lines in Fig. 1c) might be more effective if the difference between men’s and women’s perceived recognition decreases for higher amounts of received recognition.
@@
To test these relationships, we evaluate data collected from introductory physics students via an online survey, using existing measures of perceived [19, 8, 10, 11, 12] and received [13, 14, 15, 16, 17] peer recognition. The perceived recognition item asks students the extent to which they agree with the statement “My peers see me as a physics person” on a 5-point Likert scale ranging from “Strongly disagree” to “Strongly agree.” The received recognition item asks students to nominate peers who they feel are “particularly strong” in their physics course, through which we determine the number of nominations received by each student. We ask these questions separately for the instructional contexts of laboratory (lab) and lecture, when applicable, because prior work has demonstrated that peer recognition forms differently in these two instructional contexts given their different course structure and learning objectives [16, 17].
@@
We perform our analysis with 1,688 unique students across 27 introductory physics contexts (hereon referred to as individual “courses”): 15 instructional lecture courses (containing 1,039 total students, 559 men and 480 women) and 12 instructional lab courses (containing 1,652 total students, 794 men and 858 women; see Table 1). The courses come from eight PhD-granting institutions in the United States, including one Historically Black College or University and one Hispanic-Serving Institution. Four of the institutions are public and four are private. All 27 courses had a survey response rate above 70% of enrolled students. We exclude data collected from an additional 47 introductory physics courses (34 instructional lecture courses and 13 instructional lab courses, all from the same set of eight institutions) with a response rate below 70% because response rates below this threshold render our measure of received recognition too sensitive to missing data [21]. We also only include in the analysis students who self-identify as a man or a woman because prior work predicts recognition differences between these two groups–there are not yet any established expectations of recognition differences between other gender identities.
@@
We note that we examine students in physics courses pursuing a variety of majors and that such students may have different motivations for taking physics. Physics identity, however, is associated with students’ success in their physics courses, and thus is valuable for any student who is required to take physics. Moreover, physics identity has been shown to be important for the career choices of engineering students, who comprise a large fraction of our sample [22]. Non-physics students’ experiences in physics courses can also shape their other disciplinary identities and intentions to pursue their chosen field of study [23].
@@ RESULTS
Contrary to some prior work [16], we find that men and women receive proportional recognition in the lab context: the fractions of nominations made to men and women do not significantly differ from the fractions of men and women in the sample (the error bars of the proportions in lab in Fig. 2b overlap with zero; see distributions in Figs. 3b and 5). The same analysis shows that, on aggregate, men (women) receive disproportionately more (less) recognition than expected (i.e., the fraction of nominations received by men is much larger than the the fraction of men in the sample) in the lecture context (the error bars of the proportions in lecture in Fig. 2b do not overlap with zero), with a small effect size.
@@
We find that the second relationship (Fig. 1b) best describes our data set for both instructional contexts (Figs. 2c and 2d). That is, for students receiving the same amount of peer recognition (and having the same academic year, academic major, and race or ethnicity), women report significantly lower perceived recognition than men (lab: b = –0.24 [95% confidence interval: –0.35, –0.12], SE = 0.06, t = –3.95, p $<$ 0.001; lecture: b = –0.30 [–0.44, –0.15], SE = 0.08, t = –4.00, p $<$ 0.001). Received recognition also separately predicts perceived recognition: students receiving more recognition (and having the same gender, academic year, academic major, and race or ethnicity) report significantly higher perceptions of their recognition in both contexts (lab: b = 0.07 [0.01, 0.14], SE = 0.03, t = 2.27, p $<$ 0.01; lecture: b = 0.11 [0.05, 0.17], SE = 0.03, t = 3.58, p $<$ 0.001).
Comparing the sizes of these coefficient estimates, gender has a larger effect than received recognition on perceived recognition. To achieve an increase in perceived recognition equal to the average difference between men’s and women’s perceived recognition, a student would need three more nominations (i.e., a three unit increase in received recognition) in either context. Additionally, the difference between men’s and women’s perceived recognition does not vary for different amounts of received recognition as indicated by the interaction term (lab: b = –0.01 [–0.10, 0.07], SE = 0.04, t = –0.27, p = 0.79; lecture: b = –0.01 [–0.13,0.10], SE = 0.06, t = –0.23, p = 0.82). Visually, Fig. 2c indicates these results through the parallel but systematically offset lines for men and women, similar to Fig. 1b. One caveat is that we find no aggregate gender bias in received recognition in the lab context, indicating that in this context the lines for men and women actually span a comparable horizontal range (in contrast to what is shown in Fig. 1b).
@@ DISCUSSION
Results indicate that students’ varying internalization of their peer recognition better explains gender differences in student perceptions of their peer recognition than gender biases in received peer recognition: we identify that the effect of gender is larger than the effect of received recognition on perceived recognition. For men and women receiving the same peer recognition (and having the same gender, academic year, academic major, and race or ethnicity), women, on average, report their perceived peer recognition as 0.25 points (on a 5-point Likert scale) lower than men. A woman would need to receive three additional nominations from peers (received recognition) to increase their perceived peer recognition by this amount.
@@
If the goal is for men and women to hold comparable perceptions of their peer recognition, instructional interventions must aim to vertically shift men’s and women’s lines in Fig. 2c to be overlapping. In both lab and lecture contexts, this suggests that interventions should target shifts in perceived peer recognition, such that all physics students appropriately internalize recognition from their peers. Shifts in received recognition (that is, horizontal shifts in men’s and women’s lines) would not produce overlapping lines in perceived recognition.
@@
Notably, this recommendation is in contrast to existing recommendations that instructional interventions should target received recognition to mitigate perceived recognition gaps [24, 25, 16]. Our statistical results indicate that such an intervention may be particularly ineffective in an instructional lab context, where men and women receive proportional recognition from their peers as strong in the course material. Our results do, however, indicate that an intervention targeting received peer recognition (i.e., aiming to horizontally shift men’s and women’s lines in Fig. 2c to span comparable ranges) could effectively supplement an intervention targeting perceived recognition in a lecture context, where women receive significantly less peer recognition than men. That is, interventions that shift the lines in Fig. 1b vertically would result in the situation represented by Fig. 1a, requiring additional horizontal shifts for men’s and women’s recognition to be overlapping.
@@
While our findings provide strong evidence for designing interventions targeting students’ perceived recognition (and, in the lecture context, supplemental interventions targeting received recognition), our data do not allow us to make specific recommendations for what such interventions may look like. Future research will need to identify causal mechanisms of our correlational findings, answering the question: why do men and women internalize their peer recognition differently? This research will likely require both in-depth qualitative methods that identify possible mechanisms (e.g., using observations and student interviews) and controlled experiments that isolate and test individual mechanisms. Here we discuss a few plausible mechanisms to motivate this future work.
@@
These and similar lines of research will further illuminate how students internalize their peer recognition and why men and women internalize their peer recognition differently, as we observe in the large-scale analysis presented here. With these mechanisms, the research community can then design, implement, and evaluate instructional interventions that are effective in mitigating gender differences in perceived peer recognition. Such interventions will help more women and students from marginalized groups develop their physics identity and better succeed in their physics courses.

### v4

@@ Abstract
Gaining recognition as a physics person by peers is an important contributor to undergraduate students’ physics identity and their success in physics courses. Previous research has separately demonstrated that women perceive less recognition from peers than men in their physics courses (perceived recognition) and that women receive fewer nominations from their peers as strong in their physics course than men (received recognition). The relationship between perceived and received peer recognition for men and women, however, is not well understood. Here we test three plausible models for this relationship. We conduct a large-scale, quantitative study of over 1,700 students enrolled in introductory physics courses at eight institutions in the United States. We directly compare student gender, perceived recognition, and received recognition, controlling for other student demographics and course-level variability. Results show with high precision that, for students receiving the same amount of recognition, and having the same race or ethnicity, academic year, and major, women report significantly lower perceived recognition than men. These findings offer important implications for testable instructional interventions.
The under-representation of women and people of color in undergraduate physics courses and, subsequently, in physics careers is well-documented [1, 2, 3, 4]. Broadening the population of physicists would give voice to new and innovative ideas that are critical to scientific advancements [5, 6, 7]. While many structural barriers to entry and retention in physics exist for women in particular, one significant challenge is that women may both perceive [8, 9, 10, 11, 12] and receive [13, 14, 15, 16, 17] less recognition from their physics peers about their abilities as physicists than men. Perceived recognition from peers directly correlates with students’ development of their physics identity (i.e., the extent to which they believe they are a “physics person” [18]) and their persistence in physics courses and careers [19, 9, 10, 8, 20].
Existing research in physics education, however, has only examined gender differences in students’ perceived and received peer recognition separately [13, 14, 15, 16, 17, 8, 9, 10, 11, 12]. Understanding the nuance of how students internalize the peer recognition they actually receive is crucial for identifying testable instructional interventions aimed at mitigating gaps in perceived recognition between men and women. In this study, therefore, we directly compare the extent to which students feel recognized by their peers (perceived recognition) to the number of nominations they receive from their peers for being strong in their physics course (which we assume is a proxy for received recognition), as indicated on surveys. We operate under an equity of parity (or equity of outcomes) model where the instructional goal is for men and women to report comparable levels of perceived recognition, such that men and women have similar opportunities for identity development, success in physics courses, and retention in physics [21, 22].
@@
We directly test three possible quantitative relationships between perceived and received peer recognition for men and women (Fig. 1). The first relationship proposes that both men and women internalize their peer recognition similarly. Under this relationship, men only perceive more peer recognition than women [8, 9, 10, 11, 12] because they receive more peer recognition than women (Fig. 1a), as suggested by prior work [13, 15, 16, 17]. If this relationship best describes our data and our assumption that student nominations of strong peers made on a survey are an appropriate proxy for received recognition is true, then instructional intervention experiments should target received recognition. Visually, an intervention aimed at shifting the horizontal ranges of men’s and women’s lines in Fig. 1a would be more effective than one aimed at vertically shifting men’s and women’s lines to create overlapping lines for men and women (in line with the equity of parity goal). To target received recognition, instructors may try to refine students’ criteria through which they recognize their physics peers or diversify students’ peer networks in order to alleviate any gender biases in students’ nominations of strong peers.
@@
The second relationship, in contrast, is that men and women internalize their peer recognition differently. Under this relationship, men perceive more peer recognition than women [8, 9, 10, 11, 12] because they have higher perceptions of their peer recognition than women receiving the same amount of recognition (though men may also receive more peer recognition than women; Fig. 1b). If this relationship best describes our data (and our assumption that student nominations of strong peers made on a survey are an appropriate proxy for received recognition is true), instructional intervention experiments should target perceived recognition to ensure that all students appropriately internalize recognition from their physics peers. Visually, an effective intervention would aim to vertically shift men’s and women’s lines in Fig. 1b to be overlapping (a supplemental intervention aimed at shifting the horizontal ranges of men’s and women’s lines in Fig. 1b may also be appropriate if there is also a gender bias in received recognition). To target perceived recognition, instructors may help students identify different forms of recognition from others and/or give students time to reflect about the recognition they have received from others.
@@
The third relationship is that men and women internalize their peer recognition at different rates. Under this relationship, the difference between men’s and women’s perceived recognition is different at different amounts of received recognition (and men may also receive more peer recognition than women; Fig. 1c). If this relationship best describes our data (and our assumption that student nominations of strong peers made on a survey are an appropriate proxy for received recognition is true), the appropriate instructional intervention experiment would depend on the particular relationship. An intervention should target perceived recognition (i.e., vertically shifting men’s and women’s lines in Fig. 1c), for example, if the difference between men’s and women’s perceived recognition increases for higher amounts of received recognition, as shown in Fig. 1c. On the other hand, an intervention targeting received recognition (i.e., horizontally shifting men’s and women’s lines in Fig. 1c) might be more effective if the difference between men’s and women’s perceived recognition decreases for higher amounts of received recognition.
@@
To test these relationships, we evaluate data collected from introductory physics students via an online survey, using existing measures of perceived [19, 8, 10, 11, 12] and received [13, 14, 15, 16, 17] peer recognition. The perceived recognition item asks students the extent to which they agree with the statement “My peers see me as a physics person” on a 5-point Likert scale ranging from “Strongly disagree” to “Strongly agree.” The received recognition item asks students to nominate peers who they feel are “particularly strong” in their physics course, through which we determine the number of nominations received by each student. We ask these questions separately for the instructional contexts of laboratory (lab) and lecture, when applicable, because prior work has demonstrated that peer recognition forms differently in these two instructional contexts–even among the same set of students–given their different course structure and learning objectives [16].
We perform our analysis with 1,721 unique students across 27 introductory physics contexts (hereon referred to as individual “courses”): 15 instructional lecture courses (containing 1,059 total students; 559 men, 480 women, and 20 students of additional gender identities) and 12 instructional lab courses (containing 1,683 total students; 794 men, 858 women, and 31 students of additional gender identities; see Table 1). The courses come from eight institutions in the United States. These institutions are diverse across multiple dimensions, including minority-serving status, public and private institutions, and geographic regions (see Methods). All institutions, however, are PhD-granting and in the United States.
@@
All 27 courses had a survey response rate above 70% of enrolled students. We exclude data collected from an additional 47 introductory physics courses (34 instructional lecture courses and 13 instructional lab courses, all from the same set of eight institutions) with a response rate below 70% because response rates below this threshold render our measure of received recognition too sensitive to missing data [23]. We report the results for students who self-identify as a man or a woman because there were not enough students of additional gender identities to make meaningful statistical claims about them. We do, however, include these students in our statistical models and discuss their preliminary results in the Supplementary Information.
@@
We note that we examine students pursuing a variety of majors within their physics courses. Physics identity, however, is associated with students’ success in their physics courses, and thus is valuable for any student who is required to take physics. Moreover, physics identity has been shown to be important for the career choices of engineering students, who comprise a large fraction of our sample [24]. Non-physics majors’ experiences in physics courses can also shape their other disciplinary identities and intentions to pursue their chosen field of study [25].
@@ RESULTS
Contrary to some prior work [16], we find that men and women receive proportional recognition in the lab context: the fractions of nominations made to men and women do not significantly differ from the fractions of men and women in the sample. Visually, the error bars of the proportions in lab in Fig. 2b overlap with zero (see full distributions of received recognition in Figs. 3b and 5). The same analysis shows that, on aggregate, men (women) receive disproportionately more (less) recognition than expected in the lecture context: the fraction of nominations received by men is much larger than the fraction of men in the sample, with a small effect size. Visually, the error bars of the proportions in lecture in Fig. 2b do not overlap with zero.
@@
We find that the second relationship (Fig. 1b) best describes our data set for both instructional contexts (Figs. 2c and 2d). That is, for students receiving the same amount of peer recognition (and having the same academic year, academic major, and race or ethnicity), women report significantly lower perceived recognition than men (lab: b = –0.24 [95% confidence interval: –0.35, –0.12], SE = 0.06, t = –3.92, p $<$ 0.001; lecture: b = –0.30 [–0.45, –0.15], SE = 0.08, t = –4.03, p $<$ 0.001). Received recognition also separately predicts perceived recognition: students receiving more recognition (and having the same gender, academic year, academic major, and race or ethnicity) report significantly higher perceptions of their recognition in both contexts (lab: b = 0.07 [0.01, 0.14], SE = 0.03, t = 2.25, p $<$ 0.05; lecture: b = 0.11 [0.05, 0.17], SE = 0.03, t = 3.52, p $<$ 0.001).
Comparing the sizes of these coefficient estimates, gender has a larger effect than received recognition on perceived recognition. To achieve an increase in perceived recognition equal to the average difference between men’s and women’s perceived recognition, a student would need three more nominations (i.e., a three unit increase in received recognition) in either context. Additionally, the difference between men’s and women’s perceived recognition does not vary for different amounts of received recognition as indicated by the interaction term (lab: b = –0.01 [–0.10, 0.07], SE = 0.04, t = –0.26, p = 0.80; lecture: b = –0.01 [–0.13,0.10], SE = 0.06, t = –0.21, p = 0.84). Visually, Fig. 2c indicates these results through the parallel but systematically offset lines for men and women, similar to Fig. 1b. One caveat is that we find no aggregate gender bias in received recognition in the lab context, indicating that in this context the lines for men and women actually span a comparable horizontal range (in contrast to what is shown in Fig. 1b).
@@ DISCUSSION
Results indicate that, across eight different institutions, students’ varying internalization of their peer recognition better explains gender differences in student perceptions of their peer recognition than gender biases in received peer recognition. We identify that the effect of gender is larger than the effect of received recognition on perceived recognition for our sample of students. For men and women receiving the same peer recognition (and having the same gender, academic year, academic major, and race or ethnicity), women, on average, report their perceived peer recognition as 0.25 points (on a 5-point Likert scale) lower than men. A woman would need to receive three additional nominations from peers (received recognition) to increase their perceived peer recognition by this amount.
@@
For men and women to hold comparable perceptions of their peer recognition, therefore, instructional intervention experiments must aim to vertically shift men’s and women’s lines in Fig. 2c to be overlapping. In both lab and lecture contexts, this suggests that intervention experiments should target shifts in perceived peer recognition, such that all physics students appropriately internalize recognition from their peers. Shifts in received recognition (that is, horizontal shifts in men’s and women’s lines) would not produce overlapping lines in perceived recognition.
@@
Notably, this recommendation is in contrast to existing recommendations related to both peer and instructor recognition, which suggest instructional intervention experiments should target received recognition to mitigate perceived recognition gaps [26, 27, 16]. Our statistical results indicate that such an intervention may be particularly ineffective in an instructional lab context, where men and women receive proportional recognition from their peers as strong in the course material. Our results do, however, indicate that an intervention targeting received peer recognition (i.e., aiming to horizontally shift men’s and women’s lines in Fig. 2c to span comparable ranges) could effectively supplement an intervention targeting perceived recognition in a lecture context, where women receive significantly less peer recognition than men. That is, interventions that shift the lines in Fig. 1b vertically would result in the situation represented by Fig. 1a, requiring additional horizontal shifts for men’s and women’s recognition to be overlapping.
@@
While our findings provide strong evidence for designing intervention experiments targeting students’ perceived recognition (and, in the lecture context, supplemental interventions targeting received recognition), our data do not allow us to make specific recommendations for what such interventions may look like. Future research will need to identify causal mechanisms of our correlational findings, answering the question: why do men and women internalize their peer recognition differently? This research will likely require both in-depth qualitative methods that identify possible mechanisms (e.g., using observations and student interviews) and controlled experiments (similar to those reported in Refs. [28, 29]) that test individual mechanisms. Here we discuss a few plausible mechanisms to motivate this future work.
@@
Additionally, we recommend for future research to investigate the generalizability of our results by analyzing similar data from students in other countries, at other types of institutions, and from additional gender identities. We also acknowledge that recognition from people other than peers, including family members and instructors, play an important role in students’ physics identity development [19, 18]. Future work examining received and perceived recognition from these other groups of people will build on our study to create a more complete picture of the role of recognition in physics students’ outcomes.
The proposed lines of research will further illuminate how students internalize their recognition and why men and women internalize their recognition differently, as we observe in the large-scale analysis presented here. With these mechanisms, the research community can then design, implement, and evaluate instructional intervention experiments that are effective in mitigating gender differences in perceived peer recognition. Such interventions will help more women and students from other marginalized groups develop their physics identity and better succeed in their physics courses.